OpenStudy (ashy98):

Evaluate [(51 + 3) - 32] ÷ 9 • 2 A.) 52 B.) 13.3 C.) 2.5 D.) 10

OpenStudy (cj49):

54-32=22/9.2=2.5

OpenStudy (ashy98):

thank you @(^_^)@

OpenStudy (cj49):

lol...any time...